OpenStudy (anonymous):

Evaluate the logarithmic expression. Do not round the expression until the final answer. When necessary, round to the nearest hundredth. log2 32

OpenStudy (campbell_st):

ok... so this is a base 2 log so can you write 32 in terms of base of 2 and a power...?

OpenStudy (anonymous):

i think so

OpenStudy (campbell_st):

what would it be if 32 is written with a base of 2 and a power..?

OpenStudy (anonymous):

idk thats what i need help with

OpenStudy (campbell_st):

ok... this may be handy in your future math powers of 2 \[2^0 = 1\] \[2^1 = 2\] \[2^2 = 4\] \[2^3 = 8\] \[2^4 = 16\] \[2^5 = 32\] the pattern continues... but I hope you see whats happening... now a law of logs says when the base of the log and the base of the number are the same... the power is the answer... or something like that \[\log_{a} a^b = b\] so in your question you have \[\log_{2} (2^5) = ?\] any thoughts on what the answer may be..?

OpenStudy (anonymous):

is it 5

OpenStudy (campbell_st):

thats it... well done

OpenStudy (anonymous):

thanks